OpenStudy (anonymous):

f the radius of a circle is multiplied by 4, then how many times larger is the circumference of the new circle? 2. 4. 8. Pi. f the radius of a circle is multiplied by 4, then how many times larger is the circumference of the new circle? 2. 4. 8. Pi. @Mathematics

OpenStudy (anonymous):

4 times, since circumference is linearly related to circumference

OpenStudy (turingtest):

\[R_2=4R_1\]\[C_1=2 \pi R_1\]\[C_2=2\pi R_2=2 \pi(4R_1)=8\pi R_1=4C_1\]so the ansewer is 4